a quantity of 300 mL of 0.8 M HCl is mixed with 300mL of 0.4 M Ba(OH)2 in a constant pressure calorimeter having a heat capacity of 415 J/degrees C. the initial temperature of both solutions is the same at 22.0 degrees C. Given that the specific heat of the solution is 4.184J/g degrees C the density of the solution is 1.00 g/mL, and that the heat of neutralization for the process H^+(aq)+OH^-(aq)->H2O (1) is -56.2 kJ/mol, what is the final temperature of the mixed solution?
Answer #1

2 HCl(aq) + Ba(OH)2(aq) ----> Ba(Cl)2(aq) + 2H2O(l)

net ionic equation: H^+(aq)+OH^-(aq)->H2O(l) DHrxn = -56.2 kj/mol


2 mol HCl(aq) = 1 mol Ba(OH)2(aq)

no of mol of HCl taken = M*V

                        = 0.3*0.8 = 0.24 mol

no of mol of Ba(OH)2 taken = M*V

                           = 0.3*0.4 = 0.12 mol

amount of heat liberated(q) = 0.24*56.2

                            = 13.5 kj

heat released = heat absorbed by the mixture

13.5*10^3     = m*s*DT + C*DT

m = mass of reaction mixture = 600 g

s = specific heat of reaction mixture = 4.184 j/g.c

DT1 = x - 22

c = heat capacity of mixture = 415 j/c

13.5*10^3 = 600*4.184*(x-22)+415*(x-22)

x = 26.6

x = final temperature of reaction mixture = 26.6 C